OpenStudy (anonymous):

Solve the system algebraically. Check your work. 4x + 2y = 10 3x + 2y = 6

OpenStudy (anonymous):

well v cud subtract both d equations 2 eliminate y. 4x-3x+2y-2y=10-6,x=4 substitutin x in any 1 of d equations will giv us y.4(4)+2y=10,2y=10-16,y=-3 u cud check d values by pluggin dem in2 d equations. 4(4)+2(-3) =16-6=10 3(4)+2(-3) =12-6=6

OpenStudy (anonymous):

thx